Answer following questions using information from the following tables: Standard Normal Distribution Probability0.900950.99 Threshold 164 1.96 258

cum. two-tails1.00 0.50 0.40 0.30 0.20 0.10 0.05 0.02 0.01 0.002 0.001 10.000 1.000 1376 1963 3.078 6.31427 31.82 63.66 318.31 636.62 2 0.000 0.816 061 1386 1.886 2.920 4.303 6.965 9.925 22.327 31.599 1 0 5 0.000 0.727 0.920 1156 1.476 2.015 2.571 3.365 4.032 5.893 6.869 6 0.000 0.718 0.906 1.1341.4401.943 2.447 3.143 3.707 5.208 5.959 71 0.000 0.711 0.896 1.119 1.415 1.895 2.365 2.998 3.499 4.785 5.408 8 0.000 0.706 0.889 1.108 1.397 1.860 2.306 2.896 3.355 4.501 5.041 0.000 0.7030.883 1.100 1.383 1.833 2.262 2.821 3.250 4.297 4.781 00.000 0.700 0.879 1.093 1.372 1.812 2228 2.764 3.169 4.144 4.587 12 0.000 0.695 0.873 083 1356 1.782 2.179 2.681 3.055 3.930 4318 13 0.000 0.694 0.870 1.079 350 1.771 2.160 2.650 3.012 3.852 4221 6 0.000 0.690 0.865 1071 1.337 1.746 2.120 2583 2.9213.686 4.015 171 0.000 0.689 0.863 1.069 1.333 1.740 2.110 2.567 2.898 3.646 3.965 18 0.000 0.688 0.862 1.067 1.330 1.734 2.101 2.552 2.878 3.610 3.922 9 0.000 0.688 0.861 1.066 1.328 1.729 2.093 2.539 2.861 3.579 3.883 20 0.000 0.687 0860 1064 1.325 1.725 2.086 2.52828453.552 3.850 2.1 [5 points] A researcher caught 25 wild Pidgey on UI campus and found that their average weight is 4.5lbs. Find the 95% confidence interval of the population mean weight of Pidgey. The population variance of weight is known to be σ2-1.12 2.2 [5 points] A researcher caught 16 wild Rattatas on UI campus and found that their average weight is 7.8lbs. Find 1 the 95% confidence interval of the population mean weight of Ratt atas. The population variance is unknown and the researcher estimated the sample variance to be s2-2.52 2.3 [5 points] A researcher is interested in the impact of steroids on physi cal body strength. He conducted an experim ent with 16 male runners. Prior to taking steroids, the average time for 100 meters was 13.12 seconds After taking steroids, the average time became 12.64 seconds. The sample variance of time was s1.Conduct a t-test to examine if steroids had a statistically significant impact on the average time of 100 meters: Ho : μ- 13.12 and H, :Į1# 13.12. State your conclusion and specify the significance
Answer #1

2.1) Since we know the variance of the population, we use Z-score for the confidence interval. The confidence interval is

(X-Z0.95一ー, x+Zo.95 :-) = (4.5-1.96(1. Î//V25). 4.5+1.96(1.1/V25)) = Vn (4.0688, 4.9312)

2.2) Since we don't know the variance of the population, we use t-statistic (with 16-1=15 degrees of freedom) for the confidence interval. The confidence interval is

(T-to.975.15-3X+t0 975.15(7.8-2.131(2.5/16),7.8+2.131(2.5/V16))- (6.47,9.13)

2.3) Given the hypothesis, we have to do a two-tailed t-test to check for the significance.

the t-statistics is

X-13.12 12.64 13.12 _-1.92. 1.92 1/V16 The p - value is(for 15 degrees of freedom) p0.074

The p-value is 0.074 and is greater than 0.05 but less than 0.1. Thus, if our significance level is 90% we reject our null (and say that there is statistically significant impact. Also, if our significance level is 95%, we fail to reject our null (and cannot talk about significance of the impact)